OpenStudy (anonymous):

I need some help with Algebra 2.

OpenStudy (anonymous):

Solve. 9x2 – 21x + 15 = 0.

OpenStudy (anonymous):

Is it 9x2 or 9x^2 ?

OpenStudy (anonymous):

9x² sorry...

OpenStudy (anonymous):

Here are the possible answers.

OpenStudy (anonymous):

9x^2 – 21x + 15 = 0 You can factor a 3 out of every term: 3( 3x^2 - 7x + 5)

OpenStudy (jiteshmeghwal9):

\[9x^2-21x+15=0\]Dividing 9 both sides\[x^2-\left( \dfrac{21}{9} \right)x+\left( \dfrac{15}{9} \right)=0\]\[x^2-\left( \dfrac{7}{3} \right)x+\left( \dfrac{5}{3} \right)=0\]Transferring \(\dfrac{5}{3}\) R.H.S\[x^2-\left( \dfrac{7}{3} \right)x=-\left( \dfrac{5}{3} \right)\]By completing the square\[x^2-\left( \dfrac{7}{3} \right)x+\left( \dfrac{7}{6} \right)=\left( \dfrac{7}{6} \right)-\left( \dfrac{5}{3} \right)\]\[\left( x-\dfrac{7}{3} \right)^2={-1 \over 2}\]

OpenStudy (jiteshmeghwal9):

Can u solve it further for \(x\) ?

OpenStudy (anonymous):

I think so, yes thanks! :)